OpenStudy (austinl):

The electric field in a region of space has the components \(E_y = E_z = 0\) and \(E_x = (4.30 N/(C·m))x\). Point A is on the y axis at y = 3.10 m, and point B is on the x axis at x = 2.00 m. What is the potential difference (in V) \(V_B – V_A\)?

OpenStudy (austinl):

@TuringTest Mi amigo, are you any good with this sort of stuff? It seems like it should be easy, but then I start thinking... then things go downhill after that.

OpenStudy (turingtest):

draw the pic first

OpenStudy (turingtest):



OpenStudy (turingtest):

\[\Delta V=\int\vec E\cdot d\vec s=\vec E\cdot \vec s\]

OpenStudy (turingtest):

so what's the equation for \(\vec s\) is the real question here

OpenStudy (turingtest):

*expression

OpenStudy (austinl):

Stupid question, but, what is \(\vec s\) And you just kinda beat me to the punch lol

OpenStudy (turingtest):

well, it's the vector representation of the path from A to B so how might you represent the path from A to B in terms of \(\hat i\) and \(\hat j\) ?

OpenStudy (austinl):

\(\vec s=-3.1\hat i + 2 \hat j\) Does that seem right?

OpenStudy (turingtest):

exactly now just do the dot product

OpenStudy (austinl):

\(\vec E=4.30\hat i\) Is this correct, as it is denoted x?

OpenStudy (turingtest):

argh sorry bad connection :P yes, your notation is better than theirs haha, using x for a direction confuses things eventually

OpenStudy (austinl):

If so, it would be \(\Delta V=-13.33\hat i\)

OpenStudy (turingtest):

how did you get that?

OpenStudy (turingtest):

first of all, you should *never* have a vector left over after taking the dot product, since \(\hat i\cdot\hat i=1\)

OpenStudy (turingtest):

Voltage is a scalar quantity like energy. It can be positive or negative between two points, but has no direction; only magnitude

OpenStudy (turingtest):

your expression for \(\vec s\) is right, and your expression for \(\vec E\) is right, so you just didn't take the dot product right

OpenStudy (austinl):

Sorry, this connection is killing me. I think I just forgot to drop the \(\hat i\)

OpenStudy (turingtest):

No, the magnitude it wrong, too

OpenStudy (austinl):

\(4.3\times -3.1~?\)

OpenStudy (turingtest):

write out \(\vec E\cdot\vec s\)

OpenStudy (turingtest):

which direction is 3.1 in?

OpenStudy (austinl):

\(\large{\Delta V = \vec E \cdot \vec s=4.3\hat i \cdot [2\hat i -3.1\hat j]=8.6}\) Stupid error haha

OpenStudy (turingtest):

oh I see, this is my fault, I let you get away with writing the equation for s wrong the first time yes you got it :)

OpenStudy (austinl):

I just plugged it into WileyPLUS and it said it was wrong....

OpenStudy (turingtest):

let's see...

OpenStudy (turingtest):

hm... if it's not a sig fig issue I'm pretty puzzled. This is a pretty straight-forward question

OpenStudy (turingtest):

unless by them saying \[\vec E=4.3x\]they mean\[\vec E=4.3x\hat i\]

OpenStudy (turingtest):

in which case you have to integrate, as that would mean that E changes along x...

OpenStudy (austinl):

That could be... how would I go about doing that then?

OpenStudy (austinl):

Here is what it looks like exactly.

OpenStudy (turingtest):

they probably mean the fiels is proportional to x I guess, as they did in that cube problem. To get the answer you do basically the same thing, but you have to integrate at the end\[\int\vec E\cdot d\vec s=\int(4.30x\hat i)\cdot(\hat j+\hat i)ds=\int_0^24.3xdx\]then integrate from x=0 to x=2...

OpenStudy (turingtest):

field*

OpenStudy (austinl):

I literally get the same answer.

OpenStudy (turingtest):

the details of the math are out of scope of the class, so I would use a symmetry argument instead of the "line integral" that I did oh yeah?

OpenStudy (turingtest):

oh yeah, you would, by halving and then squaring x

OpenStudy (roadjester):

@TuringTest are you any good at circuits?

OpenStudy (roadjester):

@austinL how's it going?

OpenStudy (austinl):

Wait... are we doing \(V_A-V_B\) Don't we need to do \(V_B-V_A\)?

OpenStudy (turingtest):

um... right... so we are going from higher potential to lower potential, so could the problem be a minus sign? do you have unlimited chances?

OpenStudy (austinl):

AHA!!!! :D That was the problem, we were doing it backwards. The answer is -8.6

OpenStudy (austinl):

And no, 4 tries total on all of my problems

OpenStudy (turingtest):

I see, I always forget that detail :P The idea there is that because we integrated in the direction of the E-field, we are basically moving away from a positive charge, which means we are going from a higher to lower potential. Oh damn you, Benjamin Franklin, for making the electron negative and not positive.

OpenStudy (austinl):

I have always wondered how they picked these things... make North, North when referring to magnetism and negative, negative when referring to electricity.